OpenStudy (anonymous):

tan(x−π/2)∗cos(3π/2+x)−sin^3(7π/2−x)/cos(x−π/2)∗tan(3π/2=x)

OpenStudy (anonymous):

plseee reply

OpenStudy (solomonzelman):

Cos(3π/2) = ?

OpenStudy (anonymous):

i am not getting you

OpenStudy (solomonzelman):

Can you tell me what cos(3π/2) equals to ?

OpenStudy (anonymous):

cos 3*180/2=cos270

OpenStudy (solomonzelman):

Yes ! and cos(270) equals ?

OpenStudy (solomonzelman):

\[0 \times a= 0\]

OpenStudy (anonymous):

ohhh i got your point

OpenStudy (solomonzelman):

Now, re-write the question :)

OpenStudy (anonymous):

okk thankyouu

OpenStudy (solomonzelman):

I would also say that for tan(3π/2)=sin/cos= -1/ 0 = undefined

OpenStudy (anonymous):